Find the total weight of the boxes of pecans. There are 24 fewer boxes of 2lb boxes of walnuts than 3lb boxes of pecans. The total weight of all boxes is 462lbs.
Answer by ewatrrr(24785)


Hi,
Let x represent number of 3 lb boxes of pecans
then (x-24) would represent the number of 2 lb boxes walnuts
question states:
3lb*x + 2lb(x-24) = 462lbs
5x - 48 = 462lbs
5x = 510lbs
x = 102, the number of 3 lb boxes of pecans
102*3lb = 306lb would be the total weight of the boxes of pecans
checking our answer
3*102 + 2 *78 =306 + 156 = 462 lbx